If pressure of CO2 in a container is given by P=(RT/2V-b)- a/4v^2 then mass of the gas in container is

Grade:11

1 Answers

Norman Manchu
6 years ago
Use vanderwaals-real-gas-equation:P=(\frac{RT}{2V}-b)-\frac{a}{4V^{2}}
Rearranging (P+\frac{a}{4V^{2}})(2V-b)=RT
Comparing to real gas equation we are going to get this​
Mass of Carbon dioxide would be approximately 22g